Your patients needs to be treated with Antirobe drops.  Antirobe concentration is 25mg/ml.  Please determine the appropriate doses in mg and ml for the following patients:

  1. 23 lb dog, dose is 6.5 mg/kg:  ______ mg    _______ml
  2. 7 lb cat, dose is 10 mg/kg:   _______mg    ________ml
Answer #1

1lb=0.454kg

23lb=0.454×23=10.442kg

Dose is 6.5mg per kg of body wt

So dose for dog =10.4×6.5=67.8 mg

25 mg antirobe=1 ml

67.8 mg = 1÷25×67.8=2.712 ml~2.8 ml

So the answer will be 67.8 mg 2.8 ml

Now for cat

7 lb = 0.454×7=3.178~3.2 kg

10 mg per kg body wt

So dose will be =10×3.2=32 mg

So 25 mg antirobe =1 ml

32 mg=1.28 ml~1.3 ml

So the answer is 32 mg 1.3 ml